1. When the body consumes glucose, it requires oxygen gas. This process is also known as combustion, producing carbon dioxide gas and liquid water. Calculate the amount of carbon dioxide gas produced. Balanced chemical equation: C6H1206 + 6O2 6CO2 + 6H2O




2. From the data obtained on number 1, calculate the percentage yield if you have produced 0.50 g of CO2 in a laboratory experiment




1
Expert's answer
2022-05-12T00:24:02-0400

1. "C_6H_{12}O_6 +6O_2 \rightarrow 6CO_2+ 6 H_2O"

1 mole of glucose gives 6 mole CO2

2. "M(CO_2)= 12+16x2=44 g\/mole"

"n(CO_2)=0.5\/44=0.0113 mole"

"M(C_6H_{12}O_6)=12x6+12+16x6=72+12+96=180 g\/mole"

If we have x g of glucose, then

at start it is x/180 mole of glucose

It gives 6x/180=x/30 mole of CO2.


Percentage yield is "\frac{0.0113 (100)}{x/30}=33.9/x \%"